# Diversity of low-mass planet atmospheres in the C-H-O-N-S-Cl system with interior dissolution, non-ideality, and condensation: Application to TRAPPIST-1e and sub-Neptunes

### v2

@@ Abstract
A quantitative understanding of the nature and composition of low-mass rocky (exo)planet atmospheres during their evolution is needed to interpret observations. The magma ocean stage of terrestrial- and sub-Neptune planets permits mass exchange between their interiors and atmospheres, during which the mass and speciation of the atmosphere is dictated by the planet’s volatile budget, chemical equilibria, and gas/fluid solubility in molten rock. As the atmosphere cools, it is modified by gas-phase reactions and condensation. We combine these processes into an open source Python package built using JAX called Atmodeller, and perform calculations for planet sizes and conditions analogous to TRAPPIST-1e and K2-18b. For TRAPPIST-1e-like planets, our simulations indicate that CO-dominated atmospheres are prevalent during the magma ocean stage, which, upon isochemical cooling, predominantly evolve into CO2-rich atmospheres of a few hundred bar at 280 K. Around 40% of our simulations predict the coexistence of liquid water, graphite, -sulfur, and ammonium chloride—key ingredients for surface habitability. For sub-Neptune gas dwarfs, pressures are sufficiently high (GPa) that gas fugacities deviate from ideality, thereby drastically enhancing solubilities. This buffers the total atmospheric pressure to lower values than for the ideal case. These effects conspire to produce CH4-rich sub-Neptune atmospheres for total pressures exceeding 3.5 GPa, provided H/C is 100 solar and O2 moderately reducing (3 log10 units below the iron-wüstite buffer). Otherwise, molecular hydrogen remains the predominant species at lower total pressures and/or higher H/C. For all planets at high temperature, solubility enriches C/H in the atmosphere relative to the initial composition.
@@ 1 Introduction
Whilst the discovery of exoplanets has proliferated in recent years, a significant challenge lies in unravelling how their bulk interior compositions relate to atmospheric speciation. Based on current demographics studies, the most abundant kinds of planets are those with masses in between that of Earth and Neptune and are divided into ‘super-Earths’, which are rocky planets with radii less than 1.5 R⊕, and ‘sub-Neptunes’, which are planets with radii larger than 2 R⊕ that likely have H2-rich atmospheres (Bean et al., 2021). Characterising planets with possible molten surfaces (magma oceans)—such as close-in rocky and super-Earth planets in addition to sub-Neptunes with thick hydrogen (H2) envelopes—demands a comprehensive treatment of thermodynamics at the melt-atmosphere interface (Chao et al., 2021; Charnoz et al., 2023; Misener et al., 2023; Rigby et al., 2024; Seo et al., 2024; Werlen et al., 2025). This is because the mass of the melt reservoirs always predominates over that contained in the atmosphere, such that the physicochemical state of the liquid (i.e., the chemical potentials of components set by pressure, temperature and bulk composition) controls the near-surface atmospheric speciation thereby setting a lower boundary condition for the atmosphere (Hirschmann, 2012; Sossi et al., 2020). Above this boundary the atmospheric speciation may be modified by dynamics and chemistry, photochemistry, and escape (Catling & Kasting, 2017; Wordsworth & Kreidberg, 2022; Selsis et al., 2023; Krissansen-Totton et al., 2024; Owen & Schlichting, 2024; Rogers et al., 2024; Cherubim et al., 2025; Nicholls et al., 2025). Ultimately, the spectral signal in the uppermost atmosphere is most accessible for interrogation by space-based observatories such as the James Webb Space Telescope (JWST). Hence, adequately modelling physical and chemical processes at the melt-atmosphere interface is foundational to interpret current and upcoming observations.
@@
During planet formation, volatiles delivered in planetary building blocks (e.g., Alexander et al., 2012; Marty, 2012) and by capture of the nebular gas (e.g., Sharp & Olson, 2022; Young et al., 2023) likely equilibrated with the growing planet’s rocky magma ocean, whose capacity to store volatile species varies according to their solubilities. A solubility law for any given volatile species relates its dissolved concentration in magma to its fugacity, which is, in turn, related to partial pressure through a fugacity coefficient. These laws are determined either experimentally or theoretically via ab initio simulations for a set of conditions including total pressure, temperature, oxygen fugacity and melt composition that affect the stable melt and/or gas species. Solubility studies have focused on deriving relations for major atmosphere-forming species (e.g., H2O, H2, CO2, CO, CH4, N2, SO2, Cl2, and the noble gases) in silicate melts (e.g., Dixon et al., 1995; Newcombe et al., 2017; Sossi et al., 2023; Hirschmann et al., 2012; Ardia et al., 2013; Libourel et al., 2003; Boulliung & Wood, 2022; Bernadou et al., 2021; O’Neill & Mavrogenes, 2022; Thomas & Wood, 2021; Davis & Caracas, 2024; Jambon et al., 1986), and were largely conducted at moderate to high-pressure conditions (hundreds of bar to several GPa) since they were originally devised to cover the conditions of Earth’s mantle and crust. These conditions, however, are also comparable to the melt-atmosphere interface of large rocky planets and hence these solubility relations are now finding new applications in exoplanet research.
@@
For interface pressures greater than 1 GPa, gases can no longer be reasonably treated as obeying the ideal gas law, depending on temperature and the nature of the gas species (Bridgman, 1924; Duan & Zhang, 2006). Fugacity coefficients quantify how strongly a gas departs from ideal behaviour and are often expressed in terms of temperature (), pressure (), and composition () (e.g., Redlich & Kwong, 1949; Tian & Heng, 2024). Deviations from unity pertain physically to differences in volume at a given pressure and temperature relative to the ideal gas law, which are products of 1) the finite size of gas particles and 2) attractive/repulsive interactions between them (Van der Waals, 1873) and are themselves a function of , and . They can be derived from real gas equations of state (e.g., Saxena & Fei, 1988; Holland & Powell, 1991), thermodynamic tables and charts (e.g., Chase, 1998), or directly from experimental data and measurements (e.g., Presnall, 1969; Jakobsson & Oskarsson, 1994; Frost & Wood, 1997). Non-ideality is of direct relevance to interpreting the formation and evolution of rocky planets, which are mostly rock by mass and gas by volume. Both gas-phase reactions and gas solubilities are regulated by fugacities, which are only equivalent to partial pressures in the ideal limit. Meanwhile, pressure (not fugacity) is the operative quantity in satisfying volatile mass balance (e.g., Appendix C, Bower et al., 2022). In essence, non-ideality demands that ‘thermodynamic pressure’, relevant to reactions and solubility, and ‘mechanical pressure’, relevant to mass balance, are treated as distinct entities. This may lead to exoplanet demographics that are notably different from those derived by assuming ideal gas behaviour and chemically inert rocky cores (e.g., Kite et al., 2019).
@@
For planets that subsequently cool and crystallise their mantles, such as the solar system bodies, volatiles exsolve to form an early outgassed atmosphere that sets the initial state for the subsequent long-term evolution of the atmosphere (e.g., Bower et al., 2019, 2022; Nikolaou et al., 2019; Katyal et al., 2020; Lichtenberg et al., 2021; Gaillard et al., 2021; Liggins et al., 2022; Krissansen-Totton et al., 2024; Maurice et al., 2024). The gaseous species in these atmospheres, in addition to any condensed phases that form upon cooling, may be the earliest ingredients necessary for prebiotic chemical reactions to take place (Sossi et al., 2020; Zahnle et al., 2020; Benner et al., 2020; Preiner & Martin, 2021; Johansen et al., 2024; Korenaga, 2025). However, no systematic study of how volcanically degassed atmospheres cool and condense under a range of conditions distinct from that of the Earth has been undertaken. Therefore, whether the raw chemical ingredients for the origin of life were endemic to Earth, and to what extent they occur in other planetary systems, remains unanswered. Moreover, super-Earths and sub-Neptunes may sustain a molten interior, either due to close proximity to their host star or an insulating envelope, and hence volatiles could remain sequestered in the molten interior, giving rise to a wide variety of atmospheres. Hence, quantifying both the volatile budget of rocky worlds that potentially become habitable (e.g., Earth and Earth twins), as well as the interior–atmosphere exchange occurring on super-Earths and sub-Neptunes, requires compositionally-consistent and accurate volatile solubility relations applicable to the unique conditions at the interior–atmosphere interface.
@@
We have therefore devised a versatile and modular toolkit (Atmodeller) that the community can utilise and extend to compute volatile partitioning between planetary reservoirs with solubility, non-ideality, and condensation. In particular, this will enable new theoretically and experimentally-derived solubility laws and fugacity coefficients to be efficiently and accurately incorporated into models devised by (exo)planetary scientists. Ultimately, Atmodeller provides a transparent and robust framework for planetary outgassing calculations. After presenting the theory and method on which Atmodeller is based (Section 2), we then showcase its features by calculating the earliest atmosphere of TRAPPIST-1e (Section 3.1) as well as the atmosphere of a gas dwarf sub-Neptune (Section 3.2).
@@ 2.1 Formulation and numerical solution
We focus on providing an overview of the specific design considerations of Atmodeller, noting that a review of the general approaches to formulate and solve chemical equilibrium calculations are presented in Leal et al. (2017). Gas and (optionally) condensed species are specified by the user and are used to determine an independent reaction set via Gaussian elimination (see Appendix A, Leal et al., 2016). Atmodeller then applies the extended law of mass-action equations (xLMA) (Leal et al., 2016) to create a closed set of non-linear equations that combines equilibrium chemistry with elemental mass balance, cast as a root-finding problem. The xLMA has an intuitive formulation using equilibrium constants and allows for the simultaneous solution of all (user-specified) species in the system without a priori knowledge of the species’ thermodynamic stability, which is particularly important to facilitate solutions with condensates (see also Kitzmann et al., 2024). Solubility laws and real gas equations of state provide optional closure conditions, where, if desired, solubility can be neglected and the system assumed ideal. In this regard, Atmodeller can also model planetary atmospheres above chemically inert mantles or cool solidified surfaces; there is no inherent requirement for magma to be present nor the gas and condensed matter to chemically interact.
@@
If included, solubility relations ensure chemical equilibrium between a species’ fugacity and its dissolved abundance in melt (Section 2.2), whereas real gas equations of state relate a species’ partial pressure to its fugacity (Section 2.3). The mass balance is formulated by considering elemental abundances in condensates, in the molten mantle, and in the atmosphere (e.g., Appendix A, Bower et al., 2019), which requires the specification of some planetary properties (e.g., surface temperature, mantle mass, melt fraction, etc.). The mantle mass (and hence the gravitational acceleration on the planet) remains constant, irrespective of the mass of volatiles dissolved in it. However, insofar as the masses of the combined volatile elements added never exceed 3–4% of the total planetary mass, even in the case of the most volatile-rich sub-Neptunes (Section 3.2), this approximation introduces minimal error. Hence, in this formulation, the mass of volatiles in the atmosphere directly determines the magma-atmosphere surface pressure and therefore is inextricably tied to the partitioning of gaseous species between the atmosphere and molten mantle. Atmodeller allows the specification of mixed constraints, such that fugacity constraints can be combined with elemental mass constraints as long as the system of equations is closed. This facilitates a common, desirable scenario in which oxygen fugacity is defined, in absolute terms as O2, or relative to a mineralogical buffer (e.g., the iron-wüstite buffer, IW) whilst other elements have fixed abundance.
@@
Atmodeller is written in the Python programming language for seamless integration with existing modelling frameworks and leverages JAX (Bradbury et al., 2018) for high-performance computation. It consists of three subpackages that contain JAX-compliant solubility laws, real gas equations of state, and thermodynamic data, all of which can be accessed independently and used as components of Python-based modelling frameworks if desired (with or without JAX support). Notably, the automatic differentiation feature of JAX is convenient for its ability to propagate derivative information through solubility laws and real gas equations of state, which often present complex functional forms that may themselves require iterative solution during the overall solution procedure. Another feature, automatic vectorisation, enables Atmodeller to easily spawn batch calculations to efficiently explore parameter space in a single vectorised solve. Finally, just-in-time compilation enables a model to be pre-compiled, which facilitates fast Atmodeller calculations as part of a repeating workflow such as a time integration in evolution models. Atmodeller uses Optimistix (Rader et al., 2024) to solve the root-finding problem and additionally leverages Lineax for linear solves (Rader et al., 2023) and Equinox (Kidger & Garcia, 2021). As with all systems of non-linear equations, obtaining a solution requires that applied constraints (fugacity, mass) can be satisfied simultaneously with the given system of equations (mass balance, equilibrium chemistry) and closure conditions (solubility laws, free energies of individual gas species, real gas equations of state). We independently verified the results of Atmodeller by comparison to FactSage 8.2 (Bale et al., 2016) calculations assuming ideal gas behaviour (Appendix C), and the Python package has additional comparisons with condensates as well as FastChem 3.1.1 (Kitzmann et al., 2024). Furthermore, each solubility law and real gas equation of state implemented in Atmodeller includes an associated test that verifies its output against tabulated data or published figures from the original source, ensuring accuracy and reproducibility of the model components.

### v3

# Diversity of Low-mass Planet Atmospheres in the C–H–O–N–S–Cl System with Interior Dissolution, Nonideality, and Condensation: Application to TRAPPIST-1e and Sub-Neptunes

@@ Abstract
A quantitative understanding of the nature and composition of low-mass rocky (exo)planet atmospheres during their evolution is needed to interpret observations. The magma ocean stage of terrestrial- and sub-Neptune planets permits mass exchange between their interiors and atmospheres, during which the mass and speciation of the atmosphere is dictated by the planet’s volatile budget, chemical equilibria, and gas/fluid solubility in molten rock. As the atmosphere cools, it is modified by gas-phase reactions and condensation. We combine these processes into an open-source Python package built using JAX called Atmodeller, and perform calculations for planet sizes and conditions analogous to TRAPPIST-1e and K2-18b. For TRAPPIST-1e-like planets, our simulations indicate that CO-dominated atmospheres are prevalent during the magma ocean stage, which, upon isochemical cooling, predominantly evolve into CO2-rich atmospheres of a few hundred bar at 280 K. Around 40% of our simulations predict the coexistence of liquid water, graphite, -sulfur, and ammonium chloride—key ingredients for surface habitability. For sub-Neptune gas dwarfs, pressures are sufficiently high (GPa) that gas fugacities deviate from ideality, thereby drastically enhancing solubilities. This buffers the total atmospheric pressure to lower values than for the ideal case. These effects conspire to produce CH4-rich sub-Neptune atmospheres for total pressures exceeding 3.5 GPa, provided H/C is 100 solar and O2 moderately reducing (3 log10 units below the iron-wüstite buffer). Otherwise, molecular hydrogen remains the predominant species at lower total pressures and/or higher H/C. For all planets at high temperature, solubility enriches C/H in the atmosphere relative to the initial composition.
@@ 1 Introduction
Whilst the discovery of exoplanets has proliferated in recent years, a significant challenge lies in unraveling how their bulk interior compositions relate to atmospheric speciation. Based on current demographics studies, the most abundant kinds of planets are those with masses in between those of Earth and Neptune and are divided into ”super-Earths,” which are rocky planets with radii less than 1.5 , and ”sub-Neptunes,”’ which are planets with radii larger than 2 that likely have H2-rich atmospheres (Bean et al., 2021). Characterizing planets with possible molten surfaces (magma oceans)—such as close-in rocky and super-Earth planets in addition to sub-Neptunes with thick hydrogen (H2) envelopes—demands a comprehensive treatment of thermodynamics at the melt-atmosphere interface (Chao et al., 2021; Charnoz et al., 2023; Misener et al., 2023; Rigby et al., 2024; Seo et al., 2024; Werlen et al., 2025) because the mass of the melt reservoirs always predominates over that contained in the atmosphere, such that the physicochemical state of the liquid (i.e., the chemical potentials of components set by pressure, temperature, and bulk composition) controls the near-surface atmospheric speciation, thereby setting a lower boundary condition for the atmosphere (Hirschmann, 2012; Sossi et al., 2020). Above this boundary, the atmospheric speciation may be modified by dynamics, chemistry, photochemistry, and atmospheric escape (Catling & Kasting, 2017; Wordsworth & Kreidberg, 2022; Selsis et al., 2023; Krissansen-Totton et al., 2024; Owen & Schlichting, 2024; Rogers et al., 2024; Cherubim et al., 2025; Nicholls et al., 2025). Ultimately, the spectral signal in the uppermost atmosphere is most accessible for interrogation by space-based observatories such as the James Webb Space Telescope (JWST). Hence, adequately modellng physical and chemical processes at the melt-atmosphere interface is foundational to interpreting current and upcoming observations.
@@
During planet formation, volatiles delivered in planetary building blocks (e.g., Alexander et al., 2012; Marty, 2012) and by capture of the nebular gas (e.g., Sharp & Olson, 2022; Young et al., 2023) likely equilibrated with the growing planet’s rocky magma ocean, whose capacity to store volatile species varies according to their solubilities. A solubility law for any given volatile species relates its dissolved concentration in magma to its fugacity, which is, in turn, related to partial pressure through a fugacity coefficient. These laws are determined either experimentally or theoretically via ab initio simulations for a set of conditions, including total pressure, temperature, oxygen fugacity, and melt composition, that affect the stable melt and/or gas species. Solubility studies have focused on deriving relations for major atmosphere-forming species (e.g., H2O, H2, CO2, CO, CH4, N2, SO2, Cl2, and the noble gases) in silicate melts (e.g., Dixon et al., 1995; Newcombe et al., 2017; Sossi et al., 2023; Hirschmann et al., 2012; Ardia et al., 2013; Libourel et al., 2003; Boulliung & Wood, 2022; Bernadou et al., 2021; O’Neill & Mavrogenes, 2022; Thomas & Wood, 2021; Davis & Caracas, 2024; Jambon et al., 1986), and were largely conducted at moderate to high-pressure conditions (hundreds of bar to several GPa) since they were originally devised to cover the conditions of Earth’s mantle and crust. These conditions, however, are also comparable to the melt-atmosphere interface of large rocky planets and hence these solubility relations are now finding new applications in exoplanet research.
@@
For interface pressures greater than 1 GPa, gases can no longer be reasonably treated as obeying the ideal gas law, depending on temperature and the nature of the gas species (Bridgman, 1924; Duan & Zhang, 2006). Fugacity coefficients quantify how strongly a gas departs from ideal behavior and are often expressed in terms of temperature (), pressure (), and composition () (e.g., Redlich & Kwong, 1949; Tian & Heng, 2024). Deviations from unity pertain physically to differences in volume at a given pressure and temperature relative to the ideal gas law, which are products of (1) the finite size of gas particles and (2) attractive/repulsive interactions between them (Van der Waals, 1873) and are themselves a function of , and . They can be derived from real gas equations of state (EOS) (e.g., Saxena & Fei, 1988; Holland & Powell, 1991), thermodynamic tables and charts (e.g., Chase, 1998), or directly from experimental data and measurements (e.g., Presnall, 1969; Jakobsson & Oskarsson, 1994; Frost & Wood, 1997). Nonideality is of direct relevance to interpreting the formation and evolution of rocky planets, which are mostly rock by mass and gas by volume. Both gas-phase reactions and gas solubilities are regulated by fugacities, which are only equivalent to partial pressures in the ideal limit. Meanwhile, pressure (not fugacity) is the operative quantity in satisfying volatile mass balance (e.g., Appendix C, Bower et al., 2022). In essence, nonideality demands that ”thermodynamic pressure,” relevant to reactions and solubility, and ”mechanical pressure,” relevant to mass balance, are treated as distinct entities. This may lead to exoplanet demographics that are notably different from those derived by assuming ideal gas behavior and chemically inert rocky cores (e.g., Kite et al., 2019).
@@
For planets that subsequently cool and crystallize their mantles, such as the solar system bodies, volatiles exsolve to form an early outgassed atmosphere that sets the initial state for the subsequent long-term evolution of the atmosphere (e.g., Bower et al., 2019, 2022; Nikolaou et al., 2019; Katyal et al., 2020; Lichtenberg et al., 2021; Gaillard et al., 2021; Liggins et al., 2022; Krissansen-Totton et al., 2024; Maurice et al., 2024). The gaseous species in these atmospheres, in addition to any condensed phases that form upon cooling, may be the earliest ingredients necessary for prebiotic chemical reactions to take place (Sossi et al., 2020; Zahnle et al., 2020; Benner et al., 2020; Preiner & Martin, 2021; Johansen et al., 2024; Korenaga, 2025). However, no systematic study of how volcanically degassed atmospheres cool and condense under a range of conditions distinct from that of the Earth has been undertaken. Therefore, whether the raw chemical ingredients for the origin of life were endemic to Earth, and to what extent they occur in other planetary systems, remains unanswered. Moreover, super-Earths and sub-Neptunes may sustain a molten interior, either due to close proximity to their host star or an insulating envelope, and hence volatiles could remain sequestered in the molten interior, giving rise to a wide variety of atmospheres. Hence, quantifying both the volatile budget of rocky worlds that potentially become habitable (e.g., Earth and Earth twins), as well as the interior–atmosphere exchange occurring on super-Earths and sub-Neptunes, requires compositionally consistent and accurate volatile solubility relations applicable to the unique conditions at the interior–atmosphere interface.
@@
We have therefore devised a versatile and modular toolkit (Atmodeller) that the community can utilize and extend to compute volatile partitioning between planetary reservoirs with solubility, nonideality, and condensation. In particular, this will enable new theoretically and experimentally derived solubility laws and fugacity coefficients to be efficiently and accurately incorporated into models devised by (exo)planetary scientists. Ultimately, Atmodeller provides a transparent and robust framework for planetary outgassing calculations. After presenting the theory and method on which Atmodeller is based (Section 2), we then showcase its features by calculating the earliest atmosphere of TRAPPIST-1e (Section 3.1) as well as the atmosphere of a gas dwarf sub-Neptune (Section 3.2).
@@ 2.1 Formulation and numerical solution
We focus on providing an overview of the specific design considerations of Atmodeller, noting that a review of the general approaches to formulate and solve chemical equilibrium calculations are presented in Leal et al. (2017). Gas and (optionally) condensed species are specified by the user and are used to determine an independent reaction set via Gaussian elimination (see Appendix A, Leal et al., 2016). Atmodeller then applies the extended law of mass-action (xLMA) equations (Leal et al., 2016) to create a closed set of nonlinear equations that combines equilibrium chemistry with elemental mass balance, cast as a root-finding problem. The xLMA has an intuitive formulation using equilibrium constants and allows for the simultaneous solution of all (user-specified) species in the system without a priori knowledge of the species’ thermodynamic stability, which is particularly important to facilitate solutions with condensates (see also Kitzmann et al., 2024). Solubility laws and real gas EOS provide optional closure conditions, where, if desired, solubility can be neglected and the system assumed ideal. In this regard, Atmodeller can also model planetary atmospheres above chemically inert mantles or cool solidified surfaces; there is no inherent requirement for magma to be present nor the gas and condensed matter to chemically interact.
@@
If included, solubility relations ensure chemical equilibrium between a species’ fugacity and its dissolved abundance in melt (Section 2.2), whereas real gas EOS relate a species’ partial pressure to its fugacity (Section 2.3). The mass balance is formulated by considering elemental abundances in condensates, in the molten mantle, and in the atmosphere (e.g., Appendix A, Bower et al., 2019), which requires the specification of some planetary properties (e.g., surface temperature, mantle mass, melt fraction, etc.). The mantle mass (and hence the gravitational acceleration on the planet) remains constant, irrespective of the mass of volatiles dissolved in it. However, insofar as the masses of the combined volatile elements added never exceed 3%–4% of the total planetary mass, even in the case of the most volatile-rich sub-Neptunes (Section 3.2), this approximation introduces minimal error. Hence, in this formulation, the mass of volatiles in the atmosphere directly determines the magma-atmosphere surface pressure, and therefore is inextricably tied to the partitioning of gaseous species between the atmosphere and molten mantle. Atmodeller allows the specification of mixed constraints, such that fugacity constraints can be combined with elemental mass constraints as long as the system of equations is closed. This facilitates a common, desirable scenario in which oxygen fugacity is defined, in absolute terms as O2, or relative to a mineralogical buffer (e.g., the iron-wüstite buffer; IW) while other elements have fixed abundance.
@@
Atmodeller is written in the Python programming language for seamless integration with existing modeling frameworks and leverages JAX (Bradbury et al., 2018) for high-performance computation. It consists of three subpackages that contain JAX-compliant solubility laws, real gas EOS, and thermodynamic data, all of which can be accessed independently and used as components of Python-based modeling frameworks if desired (with or without JAX support). Notably, the automatic differentiation feature of JAX is convenient for its ability to propagate derivative information through solubility laws and real gas EOS, which often present complex functional forms that may themselves require iterative solution during the overall solution procedure. Another feature, automatic vectorization, enables Atmodeller to easily spawn batch calculations to efficiently explore parameter space in a single vectorized solve. Finally, just-in-time compilation enables a model to be pre-compiled, which facilitates fast Atmodeller calculations as part of a repeating workflow such as a time integration in evolution models. Atmodeller uses Optimistix (Rader et al., 2024) to solve the root-finding problem and additionally leverages Lineax for linear solves (Rader et al., 2023) and Equinox (Kidger & Garcia, 2021). As with all systems of nonlinear equations, obtaining a solution requires that applied constraints (fugacity, mass) can be satisfied simultaneously with the given system of equations (mass balance, equilibrium chemistry) and closure conditions (solubility laws, free energies of individual gas species, real gas EOS). We independently verified the results of Atmodeller by comparison to FactSage 8.2 (Bale et al., 2016) calculations assuming ideal gas behavior (Appendix C), and the Python package has additional comparisons with condensates as well as FastChem 3.1.1 (Kitzmann et al., 2024). Furthermore, each solubility law and real gas EOS implemented in Atmodeller includes an associated test that verifies its output against tabulated data or published figures from the original source, ensuring the accuracy and reproducibility of the model components.